# Curriculum-Guided Layer Scaling for Language Model Pretraining

### v2

@@ 1 Introduction
Large language models (LLMs) are typically pretrained in a single, continuous pass, processing all tokens with a uniform amount of computation regardless of their complexity or relevance to downstream tasks of interest. While this approach has shown remarkable success in large-scale models like GPT-4 (OpenAI et al., 2023) and Llama 3 (Dubey et al., 2024), it differs significantly from how humans learn, often leading to models that excel in generating coherent text but struggle with long-context reasoning across varied tasks (Schnabel et al., 2025). Recent works like Phi-3 (Abdin et al., 2024), MiniCPM (Hu et al., 2024), and others (Feng et al., 2024) have explored midtraining, adjusting the training data distribution partway through training by incorporating higher-quality, multilingual, or long-form text. However, this coarse-grained curriculum is applied on fixed model architectures. Inspired by how humans progressively build knowledge alongside their physically growing brains, we explore whether gradually scaling a model in tandem with increasingly complex data can enable more efficient and effective learning.
Curriculum learning (Bengio et al., 2009) has shown success in guiding models from easier to harder tasks, but remains largely unexplored in modern language model (LM) pretraining. Recent studies propose several alternative methods to schedule complexity and model scale during LM training. Fan & Jaggi (2023) train small proxy models to determine sample learnability, enabling efficient complexity-aware curricula for training LMs. Progressive layer stacking (Gong et al., 2019; J. Reddi et al., 2023; Saunshi et al., 2024; Gu et al., 2020; Yang et al., 2020; Panigrahi et al., 2024), layer-dropping (Zhang & He, 2020), and pruning (Kim et al., 2024) are gradual architectural adjustments that accelerate training and reduce inference cost. In computer vision, progressive model growing has been explored for image generation (Karras et al., 2017) and segmentation (Fischer et al., 2024).
In particular, progressive stacking is an approach that trains a model in stages, adding new layers (randomly initialized or copied) at the start of each stage until a final model size is eventually reached. Despite the intuitive appeal of progressive stacking as an approach to learn hierarchical features and improve compute efficiency, recent studies suggest it underperforms in knowledge-intensive tasks such as closed-book question answering. For example, MIDAS (Saunshi et al., 2024) finds that progressively grown models lag behind full-capacity models trained from scratch on memorization-focused tasks. In our own setup and compute-controlled experiments, we also find that progressive growing performs similarly or worse than standard training on several benchmarks.
@@
  • We propose Curriculum-Guided Layer Scaling (CGLS), a unified framework for scaling both model capacity and data complexity during pretraining, combining layerwise expansion and stage-wise transfer with curriculum learning.
  • We conduct a rigorous empirical evaluation of CGLS at two parameter counts and compute scales, demonstrating its scalability.
@@
    • At a LLaMA-3.2-1B parameter count and 2.5B tokens, we stratify the DataComp-LM (DCLM) corpus (Li et al., 2024) using a DistilBERT-based classifier (Sanh et al., 2019) and show that CGLS outperforms compute-matched baselines on knowledge-intensive and reasoning-heavy tasks, on average by 1.70% across all benchmarks and by as much as 5% on ARC-Easy. Scaling up to a Chinchilla-optimal 20B tokens (Hoffmann et al., 2022), gains are even larger, with an average 3.90% increase across benchmarks.
@@ 2 Related Work
Curriculum learning, formalized by Bengio et al. (2009), has shown empirical benefits in many domains, from vision (Hacohen & Weinshall, 2019; Weinshall et al., 2018) to reinforcement learning (Graves et al., 2017). However, it remains underexplored in the context of large language model (LLM) pretraining, where most models are trained in a single pass over a diverse and unstructured dataset. Self-paced learning strategies (Kumar et al., 2010) adjust training dynamics based on model uncertainty, gradually introducing harder samples as the model becomes more confident. More recently, efforts such as DataComp (Li et al., 2024), MetaCLIP (Xu et al., 2023), and Tower (Alves et al., 2024) demonstrate that large performance gains can arise from careful dataset curation, even without architectural changes. Related work like Thrush et al. (2024) and Ankner et al. (2024) investigates the use of small language models to identify high-quality data subsets that better align with LLM performance, but does not explore staged training. Likewise, Gururangan et al. (2020) show the value of domain-specific continued pretraining after generic training, indicating that structured data exposure plays a key role even in later stages. Still, these approaches do not consider whether data curricula coordinated with model growing could improve compute efficiency.
@@ 3 Method
Humans do not learn language by memorizing the full dictionary on day one. Instead, we begin with simple patterns: repetitive sounds, basic grammar, and simple visual inputs, and gradually expand to handle more abstract concepts, longer sentences, and complex reasoning as our cognitive capacities mature. This progression in data complexity is tightly coupled with the brain’s developmental growth: as new neural structures form, they scaffold increasingly sophisticated representations of the world (Kolk & Rakic, 2022). We study whether this notion of growing neural and data complexity in tandem can improve the compute efficiency of language model pretraining.
@@
This approach aims to ensure that the representations learned in prior layers are effectively utilized and not overwritten when the model size is increased. This progression is illustrated in Figure 1.
@@ 4 Experiments
The core intuition of our approach for compute-efficient pretraining is to follow a slow transition from highly structured to more diverse data while scaling model capacity. We construct pretraining curricula using the TinyStories (Eldan & Li, 2023), BookCorpus (Zhu et al., 2015), and DataComp for Language Models (DCLM) (Li et al., 2024) corpora. We first implement our approach at a GPT-2 Small scale (Radford et al., 2019), starting with a small model trained on simple, synthetic language patterns (TinyStories) and gradually increasing both data complexity and model size. Then, we scale our approach to the parameter count of Llama-3.2-1B (Dubey et al., 2024), applying it to a stratified version of the DCLM dataset, where document complexity levels (e.g., high school, undergraduate, graduate) are inferred via a DistilBERT classifier trained on a small set of GPT-4o-labeled samples.
Next, we describe this particular instantiation of data curricula that we used in our experiments.
@@ 4.1 Data Curricula
    Class Balancing & Splitting: We then upsampled any underrepresented categories, and performed a random 80/10/10 split for training, validation, and testing, respectively.
@@ 4.2 Compute-Matched Method Comparisons
  • Simple (GPT-2-Small scale only): Starting from a randomly-initialized 6 layer model, we train for one epoch on  [^1] M tokens from TinyStories. These weights are transferred to a 9 layer model, which is trained twice on 140M tokens of BookCorpus: once in the initialization phase (transferred layers frozen) and once in the full tuning phase (all layers unfrozen). We apply a similar progression at 12 layers, using 200M DCLM tokens for the initialization phase and 461M tokens for the full tuning phase.
@@
  • Curricularized baseline: This baseline simulates a coarse curriculum by training the full model sequentially on increasingly complex data subsets. For GPT-2-Small, we train on TinyStories, then BookCorpus, then DCLM. For LLaMA-3.2-1B, we follow the same staged training but use the stratified DCLM splits (high school, undergraduate, graduate).
  • Layer-scaling only: This baseline incrementally expands the model depth across stages, like CGLS, but trains on the full (unstructured) data distribution at each stage without a curriculum. This isolates the contribution of progressive stacking alone.
@@
At GPT-2-Small parameter scale, we scaled from 6 to 12 layers in 3 stages; at Llama-3.2-1B parameter scale, from 8 to 16 layers in 4 stages. We found that this starting depth yielded optimal results, as explored in the Appendix A.3. The train FLOP budgets for each stage were also optimized and allocated to favor later, larger stages (20% each for the first three stages; 40% for the final stage at Llama-3.2-1B scale), ensuring sufficient resources for fine-tuning on the most complex data. Additionally, we allocated 20% of the FLOPs for each stage to training only the new layers, and 80% for fine-tuning the full-unfrozen model. Further details on our experimental setup and ablation studies of these hyperparameters are provided in the Appendix A.1, A.3.
@@ 4.3 Evaluation
  • PIQA (Physical Interaction QA) (Bisk et al., 2019): This dataset tests physical commonsense reasoning by presenting multiple-choice questions about everyday scenarios.
@@
All tasks are evaluated in a zero-shot setting to measure base model quality without additional fine-tuning, following recent academic pretraining works such as Hwang et al. (2025), although we include some few-shot results in the Appendix A.2. We report the token-normalized accuracy, as described above, for all multiple-choice tasks. We additionally compute perplexity on a 180,000 sample subset of the Pile dataset, capturing how well the methods model diverse, noisy data from a variety of domains. We excluded LAMBADA at the GPT-2-Small scale, as our models at this compute budget perform at or near chance on this benchmark, making it uninformative for comparing pretraining strategies.
@@ 5 Results
Our experiments evaluate the performance of the CGLS training strategy at two parameter counts (matched to GPT-2-Small and Llama-3.2-1B) across perplexity and zero-shot accuracy on several downstream tasks. Table 1 summarizes the findings, comparing CGLS to several baselines with matched FLOP budgets. Across both model scales, GPT-2-Small trained on 700M tokens and Llama-3.2-1B trained on 2.5B tokens, CGLS consistently outperforms baseline approaches on reasoning-focused benchmarks such as ARC and PIQA, demonstrating its effectiveness in improving generalization under fixed compute budgets. CGLS shows particular strength in structured and knowledge-intensive tasks, supporting the benefits of aligning model growth with data complexity.
Model
Pile
Perplexity
PIQA ARC-E ARC-C HellaSwag Lambada Average
[…]
@@
The larger models demonstrate clearer trends. While perplexity results diverged from the GPT-2-Small trend, with the randomized baseline achieving the lowest perplexity across all Pile domains (see A.5 for detailed perplexity results), CGLS nonetheless achieves the strongest downstream performance. Across nearly all downstream tasks, CGLS improves over the best baseline: e.g., on PIQA (61.21% vs. 59.09%), ARC-Easy (41.37% vs. 36.36%). The gains on ARC-Easy are especially striking, exceeding the best baseline by over 5 percent. Notably, the stacking baselines, without a synchronized data curricula, did not improve downstream performance.
@@ 5.2 Additional Analyses
Beyond overall benchmark comparisons, we investigate two complementary aspects of CGLS. First, we assess stability through repeated training runs at the 2.5B-token scale (Section A.4). Across three random seeds, CGLS achieves higher performance than the baseline across all benchmarks, indicating that the method’s gains are robust to stochasticity in optimization.
@@
Model
Pile Perplexity
PIQA ARC-E ARC-C HellaSwag Lambada Average
[…]
Baseline (Randomized) 26.33 63.00% 41.79% 23.38% 33.52% 30.37% 38.41%
CGLS 22.99 67.36% 46.97% 26.28% 36.80% 34.08% 42.30%
Table 2: Evaluation results comparing CGLS and baseline approaches at the 20B token scale, aligned with the 20x data/parameter ratio in (Hoffmann et al., 2022). Both models are compute-matched, and the best model is bolded for each metric.
@@ 6 Discussion
Model
pass@8
pass@16 pass@32 pass@64
[…]
Baseline (Randomized) 3.56% 4.26% 5.27% 6.71%
CGLS 6.67% 7.63% 8.78% 9.76%
Table 3: Results for a domain-shift experiment from general-purpose web documents (DCLM) to code data (StarCoder2), using HumanEval for downstream evaluation. As a baseline, we trained a 1B parameter model on a 50/50 mix of DCLM and StarCoder2 totaling 4B tokens. CGLS was trained on precisely the same data in a different order: a curriculum gradually transitioning from general text to code, and achieves higher pass@k scores for the full range of k considered.
@@
While CGLS yields improvements across several knowledge-intensive and reasoning tasks, in a few tasks such as LAMBADA and HellaSwag, CGLS models perform comparably to the baselines, suggesting that not all tasks benefit equally from the progressive curriculum. We also observe that CGLS does not always lower perplexity: at the LLaMA-3.2-1B scale with 2.5B tokens, the randomized baseline achieved slightly better perplexity despite underperforming on reasoning tasks. Our intuition for this result is provided by studies of scaling laws with respect to downstream task performance Tay et al. (2022b; a), which find that some model inductive biases such as depth matter disproportionately more for downstream tasks than pretraining perplexity. CGLS is a curriculum over layer depth and therefore we postulate that there may be a related mechanism at play. This disconnect between perplexity and downstream performance highlights the need for better evaluation frameworks.
@@
Our experiments suggest that the initial model depth is a particularly important hyperparameter: starting with half the final number of layers ($N_{1}=N/2$) yielded the strongest performance at both pretraining scales. Starting too small (e.g., $N_{1}=6$ at LLaMA-3.2-1B scale) led to weaker final models, likely due to insufficient capacity to capture basic linguistic patterns, while starting too large (e.g., $N_{1}=10$) provided no benefit and in some cases underperformed standard training (see Section A.3). This points to a potential sweet spot in progressive scaling where initial tractability and long-term transferability are best balanced. Broadly, these results highlight a rich design space for curriculum-guided scaling. Future work could explore (1) more finely tuning the stage lengths and dataset proportions, (2) exploring other notions—e.g., syntactic diversity, lexical richness, or factuality—for defining the curriculum, (3) applying curriculum-aware optimization schedules (e.g., varying learning rates with data complexity), and (4) applying CGLS to non-text modalities (e.g., vision or multimodal data).

### v3

@@ 1 Introduction
Large language models (LLMs) are typically pretrained in a single, continuous pass, processing all tokens with a uniform amount of computation regardless of their complexity or relevance to downstream tasks of interest. While this approach has shown remarkable success in large-scale models like GPT-4 (OpenAI et al., 2023) and Llama 3 (Dubey et al., 2024), it differs significantly from how humans learn, often leading to models that excel in generating coherent text but struggle with long-context reasoning (Schnabel et al., 2025). Recent works like Phi-3 (Abdin et al., 2024), MiniCPM (Hu et al., 2024), and others (Feng et al., 2024) have explored midtraining, adjusting the training data distribution partway through training by incorporating higher-quality, multilingual, or long-form text. However, this coarse-grained curriculum is applied on fixed model architectures. Inspired by how humans progressively build knowledge alongside their physically growing brains, we explore whether gradually scaling a model in tandem with increasingly complex data can enable more efficient and effective learning.
Curriculum learning (Bengio et al., 2009) has shown success in guiding models from easier to harder tasks, but remains largely unexplored in modern language model (LM) pretraining. Recent studies propose several alternative methods to schedule complexity and model scale during LM training. Fan and Jaggi (2023) train small proxy models to determine sample learnability, enabling complexity-aware curricula for training LMs. Progressive layer stacking (Gong et al., 2019; J. Reddi et al., 2023; Saunshi et al., 2024; Gu et al., 2020; Yang et al., 2020; Panigrahi et al., 2024), layer-dropping (Zhang and He, 2020), and pruning (Kim et al., 2024) are gradual architectural adjustments that accelerate training and reduce inference cost. In computer vision, progressive model growing has been explored for image generation (Karras et al., 2017) and segmentation (Fischer et al., 2024).
In particular, progressive stacking is an approach that trains a model in stages, adding new layers (randomly initialized or copied) at the start of each stage until a final model size is reached. Despite the intuitive appeal of progressive stacking as an approach to learn hierarchical features and improve compute efficiency, recent studies suggest it under-performs in knowledge-intensive tasks such as closed-book question answering. For example, MIDAS (Saunshi et al., 2024) finds that progressively grown models lag behind full-capacity models trained from scratch on memorization-focused tasks. In our own setup and compute-controlled experiments, we also find that progressive growing performs similarly or worse than standard training on several benchmarks.
@@
  • We propose Curriculum-Guided Layer Scaling (CGLS), a unified framework for jointly scaling model capacity and data complexity via progressive layer expansion and curriculum learning.
  • We conduct a rigorous empirical evaluation of CGLS at two parameter counts and compute scales, demonstrating its scalability and robustness across both general pretraining and domain-shift settings.
@@
    • At a LLaMA-3.2-1B parameter count and 2.5B tokens, we stratify the DataComp-LM (DCLM) corpus (Li et al., 2024) using a DistilBERT-based classifier (Sanh et al., 2019) and show that CGLS outperforms compute-matched baselines on knowledge-intensive and reasoning-heavy tasks, on average by 1.70% across benchmarks and by as much as 5% on ARC-Easy. Scaling up to a Chinchilla-optimal 20B tokens (Hoffmann et al., 2022), gains are even larger, with an average 3.90% increase across benchmarks.
@@ 2 Related Work
Curriculum learning, formalized by Bengio et al. (2009), has shown empirical benefits in many domains, from vision (Hacohen and Weinshall, 2019; Weinshall et al., 2018) to reinforcement learning (Graves et al., 2017). However, it remains underexplored in the context of large language model (LLM) pretraining, where most models are trained in a single pass over a diverse and unstructured dataset. Self-paced learning strategies (Kumar et al., 2010) adjust training dynamics based on model uncertainty, gradually introducing harder samples as the model becomes more confident. More recently, efforts such as DataComp (Li et al., 2024), MetaCLIP (Xu et al., 2023), and Tower (Alves et al., 2024) demonstrate that large performance gains can arise from careful dataset curation, even without architectural changes. Related work like Thrush et al. (2024) and Ankner et al. (2024) investigates the use of small language models to identify high-quality data subsets that better align with LLM performance, but does not explore staged training. Likewise, Gururangan et al. (2020) show the value of domain-specific continued pretraining after generic training, indicating that structured data exposure plays a key role even in later stages. A closely related curriculum-learning baseline is Beyond Random Sampling (BRS) (Zhang et al., 2025), which evaluates curricula with ten difficulty groups and finds that fine-grained groupings provide little advantage over a simpler 3-tier curriculum. Still, these approaches do not consider whether data curricula coordinated with model growing could improve compute efficiency.
@@ 3 Method
Humans do not learn language by memorizing the full dictionary on day one. Instead, we begin with simple patterns: repetitive sounds, basic grammar, and simple visual inputs, and gradually expand to handle more abstract concepts, longer sentences, and complex reasoning as our cognitive capacities mature. This progression in data complexity is tightly coupled with the brain’s developmental growth: as new neural structures form, they scaffold increasingly sophisticated representations of the world (Kolk and Rakic, 2022). We study whether this notion of growing neural and data complexity in tandem can improve the compute efficiency of language model pretraining.
@@
This approach aims to ensure that the representations learned in prior layers are effectively utilized and not overwritten when the model size is increased. This progression is illustrated in Figure 1. In practice, we find that CGLS performs well with a simple default configuration across model scales: initializing with half of the final depth ($N_{1}=N/2$), allocating more compute to later stages (e.g., 20/20/20/40%), and dedicating roughly 20% of each stage to new-layer initialization. We use this configuration throughout most experiments.
@@ 4 Experiments
The core intuition of our approach for compute-efficient pretraining is to follow a slow transition from highly structured to more diverse data while scaling model capacity. We construct pretraining curricula using the TinyStories (Eldan and Li, 2023), BookCorpus (Zhu et al., 2015), and DataComp for Language Models (DCLM) (Li et al., 2024) corpora. We first implement our approach at a GPT-2 Small scale (Radford et al., 2019), starting with a small model trained on simple, synthetic language patterns (TinyStories) and gradually increasing both data complexity and model size. Then, we scale our approach to the parameter count of Llama-3.2-1B (Dubey et al., 2024), applying it to a stratified version of the DCLM dataset, where document complexity levels (e.g., high school, undergraduate, graduate) are inferred via a DistilBERT classifier trained on a small set of GPT-4o-labeled samples. Next, we describe this particular instantiation of data curricula that we used in our experiments.
@@ 4.1 Data Curricula
    Class Balancing & Splitting: We then upsampled any underrepresented categories, and performed a random 80/10/10 split for training, validation, and testing.
@@ 4.2 Compute-Matched Method Comparisons
  • Simple (GPT-2-Small scale only): Starting from a randomly-initialized 6 layer model, we train for one epoch on  [^1] M tokens from TinyStories. These weights are transferred to a 9 layer model, which is trained twice on 140M tokens of BookCorpus: once in the initialization phase (transferred layers frozen) and once in the full tuning phase (all layers unfrozen). We apply a similar progression at 12 layers, using 200M DCLM tokens for the first phase and 461M tokens for the full tuning phase.
@@
  • Curricularized baseline: This baseline simulates a coarse curriculum by training the full model sequentially on increasingly complex data subsets. For GPT-2-Small, we train on TinyStories, then BookCorpus, then DCLM. For LLaMA-3.2-1B, we follow the same staged training but use the stratified DCLM splits (high school, undergraduate, graduate). Both this curriculurized baseline and CGLS use the same difficulty classifier, the same three difficulty tiers, and the same pacing schedule; all optimization hyperparameters and compute budgets are matched. The only difference is whether the model depth is expanded during training.
  • Layer-scaling only: This baseline incrementally expands the model depth across stages, in the same schedule and manner as CGLS, but trains on the full (unstructured) data distribution at each stage without a curriculum. This isolates the contribution of progressive stacking alone.
@@
At GPT-2-Small parameter scale, we scaled from 6 to 12 layers in 3 stages; at Llama-3.2-1B parameter scale, from 8 to 16 layers in 4 stages. As explored in the Appendix A.3, initializing with half of the final model depth, and allocating more training budget to the larger states (e.g. 20% each for the first three stages; 40% for the final stage) consistently yielded strong performance and serves as a reliable default. Further details on our experimental setup are provided in the Appendix A.1.
@@ 4.3 Evaluation
  • PIQA (Physical Interaction QA) (Bisk et al., 2019): This dataset tests physical commonsense reasoning using multiple-choice questions about everyday scenarios.
@@
At our largest experiment scale of 1B parameters trained on 20B tokens, we additionally test the following benchmarks:
  • GPQA (Graduate-Level Physics QA) (Rein et al., 2023): A challenging benchmark designed to assess high-level scientific and mathematical reasoning. Questions are drawn from graduate physics coursework and require multi-step conceptual understanding.
  • TLDR9+ (Summarization) (Sotudeh et al., 2021): A summarization task based on Reddit posts, requiring abstractive generation of concise summaries from long, noisy inputs. We report Rouge-L (Lin, 2004) following prior work.
  • OpenRewriteEval (ORE) (Shu et al., 2023): An open-ended rewriting and paraphrasing benchmark. The model must rewrite user-provided text to satisfy semantic constraints such as preserving meaning, altering tone, or improving clarity. We report Rouge-L (Lin, 2004) here as well.
  • InfiniteBench EN.QA (Long-Context QA) (Zhang et al., 2024): A long-context benchmark where models must answer questions requiring retrieval and reasoning over very lengthy contexts. We report the English QA subset (EN.QA), which stresses information retention and retrieval at sequence lengths far beyond typical context windows. We report ROUGE F1 scores as specified for this benchmark.
All tasks are evaluated in a zero-shot setting to measure base model quality without additional fine-tuning, following recent academic pretraining works such as (Hwang et al., 2025), although we include few-shot results in A.2. We report the token-normalized accuracy, as described above, for all multiple-choice tasks. We additionally compute perplexity on a 180,000 sample subset of the Pile dataset, capturing how well the methods model diverse, noisy data from a variety of domains. We excluded LAMBADA at the GPT-2-Small scale, as our models at this compute budget perform at or near chance on this benchmark, making it uninformative for comparing pretraining strategies.
@@ 5 Results
Our experiments evaluate the performance of CGLS at two parameter counts (matched to GPT-2-Small and Llama-3.2-1B) across perplexity and zero-shot accuracy on several downstream tasks. Table 1 summarizes the findings, comparing CGLS to several baselines with matched FLOP budgets. Across both model scales, GPT-2-Small trained on 700M tokens and Llama-3.2-1B trained on 2.5B tokens, CGLS consistently outperforms baseline approaches on reasoning-focused benchmarks such as ARC and PIQA, demonstrating its effectiveness in improving generalization under fixed compute budgets. CGLS shows particular strength in structured and knowledge-intensive tasks, supporting the benefits of aligning model growth with data complexity.
Model The Pile PIQA ARC-E ARC-C HellaSwag Lambada Average
GPT-2-Small
@@
The larger models demonstrate clearer trends. While perplexity results diverged from the GPT-2-Small trend, with the randomized baseline achieving the lowest perplexity across all Pile domains (see A.6 for detailed perplexity results), CGLS nonetheless achieves the strongest downstream performance. Across nearly all downstream tasks, CGLS improves over the best baseline: e.g., on PIQA (61.21% vs. 59.09%), ARC-Easy (41.37% vs. 36.36%). The gains on ARC-Easy are especially striking, exceeding the best baseline by over 5 percent. Notably, the stacking baselines, without a synchronized data curricula, did not improve downstream performance.
@@ 5.1 Extending CGLS
Model
pass@8
pass@16 pass@32 pass@64
Llama-3.2-1B
Baseline (Rand.) 3.56% 4.26% 5.27% 6.71%
CGLS 6.67% 7.63% 8.78% 9.76%
Table 2: Results for a domain-shift experiment from general-purpose web documents (DCLM) to code data (StarCoder2), using HumanEval for downstream evaluation. As a baseline, we trained a 1B parameter model on a randomized 50/50 mix of DCLM and StarCoder2 totaling 4B tokens. CGLS was trained on precisely the same data in a different order: a curriculum gradually transitioning from general text to code, and achieves higher pass@k scores for the full range of k considered.
@@
Model Pile PIQA ARC-E ARC-C HS LBD Avg GPQA T9+ ORE IB
Llama-3.2-1B (20B Tokens)
BL-R 26.33 63.00% 41.79% 23.38% 33.52% 30.37% 38.41% 25.22% 2.787 17.30% 0.235%
BL-C 24.58 63.55% 42.97% 24.83% 34.61% 32.66% 39.72% 22.32% 2.947 20.61% 0.401%
CGLS 22.99 67.36% 46.97% 26.28% 36.80% 34.08% 42.30% 29.24% 3.568 17.03% 0.539%
Table 3: Evaluation results comparing CGLS and baseline approaches (BL-R refers to the compute-matched baseline trained on randomized data, and BL-C refers to the compute-matched baseline trained with a curriculum) at the 20B-token scale, aligned with the 20× data/parameter ratio in (Hoffmann et al., 2022). All models are compute-matched. The listed benchmarks are: perplexity on The Pile (Pile), PIQA (PIQA), ARC-Easy (ARC-E), ARC-Challenge (ARC-C), HellaSwag (HS), Lambada (LBD), the average over MC-QA tasks (Avg), GPQA (GPQA), TLDR9+ summarization (T9+), OpenRewriteEval (ORE), and the InfiniteBench long-context EN.QA task (IB). The best score for each metric is bolded, with the second-best method underlined.
@@ 5.2 Additional Analyses
Beyond overall benchmark comparisons, we investigate two complementary aspects of CGLS. First, we assess stability through repeated training runs at the 2.5B-token scale (Section A.4), and through replacing our DistilBERT-based difficulty classifier with the FineWeb educational classifier (Penedo et al., 2024) (Section A.5). Across three random seeds, CGLS achieves higher performance than the baseline across all benchmarks, indicating that the method’s gains are robust to stochasticity in optimization. When using the FineWeb classifier, CGLS attains an average accuracy nearly identical to that obtained with the DistilBERT classifier, suggesting that the method is robust to the specific curriculum signal used to stratify data.
@@ 6 Discussion
While CGLS yields improvements across several knowledge-intensive and reasoning tasks, in a few tasks such as LAMBADA and HellaSwag, CGLS models perform comparably to the baselines, suggesting that not all tasks benefit equally from the progressive curriculum. We also observe that CGLS does not always lower perplexity: at the LLaMA-3.2-1B scale with 2.5B tokens, the randomized baseline achieved slightly better perplexity despite underperforming on reasoning tasks. Our intuition for this result is provided by studies of scaling laws with respect to downstream task performance (Tay et al., 2022a, b), which find that some model inductive biases such as depth matter disproportionately more for downstream tasks than pretraining perplexity. CGLS is a curriculum over layer depth and therefore we postulate that there may be a related mechanism at play. This disconnect between perplexity and downstream performance highlights the need for better evaluation frameworks.
@@
Our experiments suggest that the initial model depth is a particularly important hyperparameter: starting with half the final number of layers ($N_{1}=N/2$) yielded the strongest performance at both pretraining scales. Starting too small (e.g., $N_{1}=6$ at LLaMA-3.2-1B scale) led to weaker final models, likely due to insufficient capacity to capture basic linguistic patterns, while starting too large (e.g., $N_{1}=10$) provided no benefit and in some cases underperformed standard training (see Section A.3). This points to a potential sweet spot in progressive scaling where initial tractability and long-term transferability are best balanced. Additionally, our experiments with the FineWeb classifier suggest that curriculum construction is an important design choice, and that exploring improved or domain-adaptive difficulty classifiers may further strengthen CGLS. Broadly, these results highlight a rich design space for curriculum-guided scaling. Future work could explore (1) more finely tuning the stage lengths and dataset proportions, (2) exploring other notions—e.g., syntactic diversity, lexical richness, or factuality—for defining the curriculum, (3) applying curriculum-aware optimization schedules (e.g., varying learning rates with data complexity), (4) extending CGLS to larger parameter regimes (e.g. 7B), and (5) applying CGLS to non-text modalities (e.g., vision or multimodal data).